OpenStudy (anonymous):

(x - 3) - 2(x + 6) = -5

OpenStudy (anonymous):

what do you need help with? Like what don;t you understand?

OpenStudy (anonymous):

don't*

OpenStudy (anonymous):

i dont understand anything

OpenStudy (anonymous):

(x-2x)-2*6=-5 (-x)-12=-5 -x=-5+12 -x=7 x=-7

OpenStudy (anonymous):

@leedomathgeek please don't give answers

OpenStudy (kamille):

(x - 3) - 2(x + 6) = -5 (x-3) -2x-12=-5 x-3-2x-12=-5 Now, put all numbers which do not have "x" to the right side. What do you get?

OpenStudy (anonymous):

Ops....just new to this awesome website

OpenStudy (anonymous):

im kinda new to but it doesn't help if you only give theh people answers

OpenStudy (anonymous):

the*